The area of triangle ABC is 20 square units. If angle A = 130o and b = 6, find the value of c.
Answer by fractalier(2101)

We use the formula
Area = (1/2)bc sin A and plug in to get
20 = (1/2)(6)c sin 130
c = 20 / 3 sin 130
c = 8.7